OpenStudy (anonymous):

verify the identity sin theta + sin 3 theta/ cos theta + cos 3 theta = tan 2 theta

OpenStudy (mertsj):

\[\frac{\sin \theta +\sin 3\theta}{\cos \theta+\cos 3\theta}=\tan 2\theta \]

OpenStudy (mertsj):

\[\sin3 \theta+\sin \theta=2\sin \frac{3\theta+\theta}{2}\cos \frac{3\theta-\theta}{2}\]

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

but when i went to simplify i nee to get rid of the 2 in front of the sin and cos

OpenStudy (mertsj):

Use that and also this: \[\cos 3\theta+\cos \theta =2\cos \frac{3\theta+\theta}{2}\cos \frac{3\theta-\theta}{2}\]

OpenStudy (mertsj):

\[\frac{2\sin 2\theta \cos \theta}{2\cos 2\theta \cos \theta}=\tan 2\]

OpenStudy (mertsj):

tan 2 theta that is

OpenStudy (mertsj):

You follow?